OpenStudy (anonymous):

What is the simplified form of the expression 8the square root of v squared. − the square root of v to the fourth power. − the square root of the quantity 9 times v squared. + the square root of the quantity 16 times v to the fourth power.

OpenStudy (imstuck):

What does this mean...8the square root of v squared.

OpenStudy (imstuck):

It's rather unclear

OpenStudy (anonymous):

:L

OpenStudy (anonymous):

hold on

OpenStudy (imstuck):

ok

OpenStudy (anonymous):

\[8\sqrt{v^2}-\sqrt{v^4}-\sqrt{v^2}+\sqrt{16v^4}\]

OpenStudy (imstuck):

ok, the v^2 comes out as just a v. The other powers...if they are even powers, like a 4 or a 6 or an 8, they come out with half the original power. The square root of v^4 is v^2. The square root of v^6 comes out as v^3. See? So what you have here is this:

OpenStudy (imstuck):

\[8v-v ^{2}-v+4v ^{2}\]And of course you can combine and simplify...\[3v ^{2}+7v\]

OpenStudy (anonymous):

can i get an answer real quick i need to go

OpenStudy (anonymous):

please

OpenStudy (imstuck):

Are you supposed to simplify further? You could factor out a v and get this: v(3v+7)

OpenStudy (imstuck):

It's either one. 3v^2+7v or simplfied is v(3v+7)

OpenStudy (anonymous):

do u need the choices

OpenStudy (imstuck):

Does one of those match yours?

OpenStudy (anonymous):

15v2 + 11v 11v2 + 15v 5v2 + 3v 3v2 + 5v

OpenStudy (anonymous):

nope

OpenStudy (imstuck):

ok you must have mixed up a sign. I got a 7v and it should probably be a 5v. I'd say it was the last one.

OpenStudy (anonymous):

k ty